# Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models

### v2

@@ 4.2 HNC
In this section, we utilize only the questions and ground truth from the PRM800K dataset [7], without relying on manually annotated data. We employ MCTS with Qwen2.5-7B-Math-Instruct [21] as an automatic annotation method to generate reasoning trajectories. As mentioned in Section 3.2, these auto-annotated reasoning trajectories from MCTS are used to train PRM, after which we apply the HNC data augmentation method to generate additional training data for HRM.
To balance computational efficiency and robustness, we configure MCTS with 5–6 child nodes per parent and a maximum tree depth of 7, ensuring reasoning completion within 7 steps. Since the computational cost of MCTS rollouts grows exponentially with tree depth and branching factor, we limit these parameters to maintain feasibility. The full MCTS simulation requires approximately 2,457 A100-80GB GPU-hours, while the HNC augmentation process takes around 30 minutes.
We perform supervised fine-tuning of Qwen2.5-1.5B-Math [21] for both PRM and HRM. To evaluate performance, we employ different policy models, including Qwen2.5-7B-Math-Instruct [21], DeepSeek-Math [17], and Qwen2.5-72B-Math-Instruct [21], applying the best-of-$N$ strategy on the PRM800K dataset. Detailed training information is provided in Appendix A.2.
@@
Table 4: Accuracy of different policy models under PRM and HRM using the best-of-$N$ strategy on the PRM800K test set [7]. The training data for both PRM and HRM are derived from MCTS with Qwen2.5-7B-Math-Instruct [21]
@@
The relatively lower performance of Qwen2.5-72B-Math-Instruct (as shown in Table 4) can be attributed to the tree height constraints imposed by MCTS, which require answer generation to follow a predefined template within a fixed number of steps. While Qwen2.5-72B-Math-Instruct exhibits strong reasoning capabilities [21], its highly structured training process makes it more vulnerable to performance degradation when deviating from its learned format.
@@ 4.3 Self-Training
We fine-tune Qwen2.5-7B-Math-Instruct [21] using high-MC-score reasoning data extracted from the PRM and HRM training datasets. Qwen-7B-HRM denotes the policy model fine-tuned on high-MC-score reasoning data from HRM’s training set, while Qwen-7B-PRM follows the same procedure for PRM’s training set. We set $\lambda$ to 0.5. Table 5 further validates that SFT enhances the policy model’s reasoning capability by leveraging high-quality data, with HRM demonstrating greater robustness compared to PRM.
@@ 4.4 HRM Generalization Across Different Domains
To broaden the applicability of HRM and evaluate its generalization capability, we assess HRM and PRM, trained on the PRM800K dataset, on the Math500 [7] and GSM8K [6] datasets. Table 6 shows that HRM exhibits greater robustness across different domains, demonstrating superior generalization performance, particularly in Math500, where it effectively handles complex mathematical reasoning tasks.
@@ A.1 HRM Training Details
To accelerate the training process of reward model, we employ FlashAttention [29, 30], DeepSpeed [31, 32], and mixed-precision training [33, 34]. However, within the PRM800K domain, we frequently encounter the issue: "Current loss scale already at minimum - cannot decrease scale anymore. Exiting run." This indicates that the numerical precision is insufficient for stable training. To mitigate this issue and ensure reproducibility, we set max_grad_norm to 0.01, which effectively stabilizes the training process.

### v3

@@ 4.2 HNC
In this section, we utilize only the questions and ground truth from the PRM800K dataset Lightman et al. (2023), without relying on manually annotated data. We adopt MCTS as the automatic annotation method, using Qwen2.5-7B-Math-Instruct Yang et al. (2024b) as the reasoning engine to generate trajectories. As mentioned in Section 3.2, these auto-annotated reasoning trajectories from MCTS are used to train PRM, after which we apply the HNC data augmentation method to generate additional training data for HRM.
To balance computational efficiency and robustness, we configure MCTS with 5–6 child nodes per parent and a maximum tree depth of 7, ensuring reasoning completion within 7 steps. Since the computational cost of MCTS rollouts grows exponentially with tree depth and branching factor, we limit these parameters to maintain feasibility. The full MCTS simulation requires approximately 2,457 A100-80GB GPU-hours, while the HNC augmentation process takes around 30 minutes with CPU.
We perform supervised fine-tuning of Qwen2.5-1.5B-Math Yang et al. (2024b) for both PRM and HRM. To evaluate performance, we employ different policy models, including Qwen2.5-7B-Math-Instruct Yang et al. (2024b), DeepSeek-Math-7B Shao et al. (2024), and Qwen2.5-72B-Math-Instruct Yang et al. (2024b), applying the best-of-$N$ strategy on the PRM800K dataset. Detailed training information is provided in Appendix A.2.
Table 4: Accuracy of different policy models under PRM and HRM using the best-of-$N$ strategy on the PRM800K test set Lightman et al. (2023). The training data for both PRM and HRM are derived from MCTS with Qwen2.5-7B-Math-Instruct Yang et al. (2024b).
@@
The relatively lower performance of Qwen2.5-72B-Math-Instruct (as shown in Table 4) can be attributed to the tree height constraints imposed by MCTS, which require answer generation to follow a predefined template within a fixed number of steps. While Qwen2.5-72B-Math-Instruct exhibits strong reasoning capabilities Yang et al. (2024b), its highly structured training process makes it more vulnerable to performance degradation when deviating from its learned format.
@@ 4.3 Self-Training
We fine-tune Qwen2.5-7B-Math-Instruct Yang et al. (2024b) using high-MC-score reasoning data extracted from the PRM and HRM training datasets. Qwen-7B-HRM denotes the policy model fine-tuned on high-MC-score reasoning data from HRM’s training set, while Qwen-7B-PRM follows the same procedure for PRM’s training set. We set $\lambda$ to 0.5. Table 5 further validates that SFT enhances the policy model’s reasoning capability by leveraging high-quality data, with HRM demonstrating greater robustness compared to PRM.
@@ 4.4 HRM Generalization Across Different Domains
Trained solely on PRM800K, HRM and PRM are further evaluated on MATH500 Lightman et al. (2023) and GSM8K Cobbe et al. (2021) to assess their generalization to out-of-domain reasoning tasks. Table 6 shows that HRM exhibits greater robustness across different domains, demonstrating superior generalization performance, particularly in MATH500, where it effectively handles complex mathematical reasoning tasks.
@@ A.1 HRM Training Details
To accelerate the training process of reward model, we employ FlashAttention Dao et al. [2022], Dao [2024], DeepSpeed Rajbhandari et al. [2020], Aminabadi et al. [2022], and mixed-precision training Kalamkar et al. [2019], Micikevicius et al. [2017]. However, within the PRM800K domain, we frequently encounter the issue: "Current loss scale already at minimum - cannot decrease scale anymore. Exiting run." This indicates that the numerical precision is insufficient for stable training. To mitigate this issue and ensure reproducibility, we set max_grad_norm to 0.01, which effectively stabilizes the training process.